OpenStudy (anonymous):

Which is the graph of f(x) = 100(0.7)x?

OpenStudy (hockeychick23):

do you have answer choices? simplified your equation is f(x)=70x which is the same thing as f(x)=70x+0. Your graph will look like this: i don't see your choices though

OpenStudy (mathmate):

I think you mean \(f(x)=100(0.7)^x\) The graph of \(f(x)=a(b^x)\) looks like this:  The graph intersects the y-axis at (0,a). When b>1, the curve rises (f(x) increases with x) When b<1, the curve falls (f(x) decreases with x) So substitute your parameters and find the right curve.